HETAV PATEL

Grade 11,

The area bounded by the x-axis and y=cosx and y=sinx when 0 x pi/2 Pls reply with the best soln as fast as possible !!! Always gonna be approved !!!

The area bounded by the x-axis and y=cosx and y=sinx when 0<x<pi/2


 


Pls reply with the best soln as fast as possible !!!


Always gonna be approved !!!

Grade:12

2 Answers

vikas askiitian expert
13 years ago

f(x) = cosx , g(x) = sinx

 

these two curves intersect at x = pi/4 ,now draw the graphs & u will observe  that this area

is divided  by x = pi/4 so one half area is under sinx & other half is under cosx so area is

 

A = A1 + A2

 

A =  A1 = (sinxdx) lim from 0 to pi/4    +    A2 = (cosxdx)lim from pi/4 to pi/2

A =  (1  -  1/root2)       +     (1 - 1/root2)

A = 2 - root2 square units

ramendra bahadur singh
13 years ago

ITS AREA =2-ROOT2 Sq.UNITS